• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ransome

    Log entry time 08:01:14 on January16,2009

    Entry number 256311

    keyword=Owl Shift summary

    Begin shift with run 5201 in progress
    5201 ended 12:32 1 million, 0.047 C
    5202 12:32 to 1:35 1 million, 0.047 C
    5203 1:35 to 2:40 1 million .047
    5204 2:40 to 3:45 1 million .047
    5205 3:45 to 4:52 1 million .047
    5206 4:53 to 6:10 1 million .047

    Raise beam current to 14 uA.
    5207 6:12 to 7:30 (beam off at 6:55) 800 K .036 C

    7:10 - MCC calls - they had a fire alarm, smoking heater.
    Beam has been off since 6:55, they expect no beam for another
    20-30 minutes.
    7:30 Beam back
    5208 7:30 - still running at end of shift

    Good shift